A box has the following dimensions:
Length 100cm
Width: 80cm
Heigh: 120cm
a) What is the Volume of the box measured in cubic meters?
b) What is the Volume of the box measured in cubic centimeters?

Found 2 solutions by Boreal, natolino_2017:
Answer by Boreal(15235)

Volume in cc is 100*80*120=960,000 cm^3
1m^3=(100cm/m)^3=100000 cm^3
Therefore, 960000cm^3=0.96 m^3
=====================
Check
1m long*0.8 m wide*1.2 cm high=0.96m^3.

Answer by natolino_2017(77)

the volume of a box is the product of the (legth)(width)(height) if they are in the same units.
V = 100*80*120 = 960,000 cm^3 = 960 liters.
@natolino_